### v2

@@ 3.3 Experimental Setup
Datasets. We evaluate our work on two publicly available surgical phase datasets: Cholec80 (C80) [32] and AutoLaparo21 (AL21) [33]. For C80, we follow the 40/40 train-test split for classification as in [20, 4], and for the regression task, we use the 60/20 train-test split following [12, 13]. For AL21, we follow the same 14/7 split for training and testing as in [4]. Both datasets were sampled at 1 frame per second (fps) following previous works [20, 4]. Across both datasets, we train on the 7 surgical phases, including an additional end-of-surgery (EOS) class, while disregarding tool annotations.
Training Details. We train a single model on a maximal horizon $H=60$ minutes and evaluate it at intermediate steps without re-training. We use a simple cross-entropy loss for the anticipation task and mean squared error loss for the regression task. Additional details on the training procedure are provided in the supplementary material (Sec. 1.1 in SM).
Anticipation Time. We assess our method over a range of intermediate horizons to capture both short-term (2, 3, and 5 minutes) and long-term (15, 30, 45, and 60 minutes) events. Given that C80 procedures have an average duration of 38 minutes and AL21 procedures average 66 minutes, extending phase predictions up to 60 minutes captures most of the workflow. This range enables us to simultaneously predict immediate transitions and estimate the remaining surgery duration.
Classification Task. Since long-term future phase classification has not been studied in previous work, direct comparisons to other surgical workflow anticipation methods are not possible. Therefore, we assess our generative approaches against two baseline methods: a simple continuation model (Naive1) and a probabilistic prediction model (Naive2). Naive1 baseline extends the current recognised class over the future predicted horizon for $H$ minutes. While effective for short-term predictions, its performance naturally degrades over time with new class transitions. Naive2 baseline predicts future phases by sampling from the probability distribution $P_{i,j,n_{i}\cdot 60}$, conditioned on the predicted current class $i$ and the anticipated future index $n_{i}$. This baseline is effective due to the strong priors inherent in structured workflows, leveraging high recognition accuracy and precisely defined anticipation times.
@@ 3.4 Evaluation Metrics
We use a combination of classification and regression metrics to evaluate our methodology. These metrics are calculated for each anticipation time, allowing us to analyse the model’s performance through time. We use frame-based accuracy to measure the proportion of correctly classified frames among all predictions. Although the videos are untrimmed with preserved class distribution from full surgeries, we still provide class-based metrics (F1-score, Precision, and Recall) over extended time horizons in the supplementary materials to assess performance on less frequent phases. Finally, we use the Mean Absolute Error (MAE) to evaluate the model’s performance on the regression task, specifically for predicting the remaining time until the next phase occurrence. This metric quantifies the average absolute prediction errors in minutes. We use the wMAE, inMAE, and outMAE for weighting samples outside and inside the temporal horizon as proposed in [12, 13], respectively. We select the best model based on the mean accuracy and wMAE over each task’s time intervals.
@@ 4.1 Surgical Phase Anticipation
Table 1 summarises the classification results for phase recognition at and anticipation at specific future time points $\tau_{5}$ and $\tau_{15}$, where $\tau_{i}=t+n_{i}\cdot 60$ (in minutes). Results are reported for both Cholec80 and AutoLaparo21 datasets.
Table 1: Surgical Phase Anticipation. $Acc_{\tau_{t}}$: phase anticipation at time $t$ (in minutes). $Acc_{\tau_{\forall}}$: mean accuracy for all anticipation time steps.
Methods Cholec80 (40/40) AutoLaparo21 (14/7)
$\mathbf{Acc_{\tau_{0}}}$ $\mathbf{Acc_{\tau_{5}}}$ $\mathbf{Acc_{\tau_{15}}}$ $\mathbf{Acc_{\tau_{\forall}}}$ $\mathbf{Acc_{\tau_{0}}}$ $\mathbf{Acc_{\tau_{5}}}$ $\mathbf{Acc_{\tau_{15}}}$ $\mathbf{Acc_{\tau_{\forall}}}$
Naive1 † 90.9 47.5 12.0 32.2 78.0 52.0 15.2 35.5
Naive2 † 90.9 49.0 44.1 47.8 78.0 45.2 35.6 41.9
SWAG-AR 90.4 56.4 29.4 44.7 75.5 51.0 21.5 38.1
SWAG-SP 89.0 43.7 46.4 45.8 75.7 51.1 44.8 48.1
SWAG-SP* 89.8 51.5 47.0 49.5 74.4 51.4 45.3 49.8
SWAG-SP-R2C 90.9 59.6 54.5 56.6 78.0 52.1 46.0 47.5
  • Note: The light blue cells in bold and underlined values are the best and second-best results, respectively. * indicates the presence of our prior knowledge token embedding initialization method. † indicates models that were only supervised on the recognition task ($Acc_{\tau_{0}}$).
On average over all horizons $Acc_{\tau_{\forall}}$, SWAG-SP-R2C achieves superior performance on Cholec80 and ranks second on AutoLaparo21. Figure 3 shows that SWAG-SP-R2C and SWAG-SP* outperform other methods and baselines on both datasets over an 18-minute horizon. SWAG-SP* shows a marked improvement in accuracy, highlighting the value of our prior knowledge embedding strategy, especially on Cholec80. The relatively strong performance of Naive2 and SWAG-SP* on Cholec80 compared to AutoLaparo21 likely reflects a more uniform probability distribution of possible future phases in the AutoLaparo21 dataset.
Figure 3: Model performance (accuracy) of surgical phase recognition and anticipation on Cholec80 and AutoLaparo21 up to 18 minutes, with mean values over 18 minutes.
@@ 4.2 Surgical Phase Anticipation with Remaining Time
Based on the stronger classification results from the single-pass approaches i.e. SWAG-SP* and SWAG-SP-R2C, we select the single-pass (SWAG-SP) decoder for comparison on state-of-the-art methods on the remaining time until the next phase occurrence regression task (see Table 2).
@@
$wMAE[\downarrow]$ $inMAE[\downarrow]$ $outMAE[\downarrow]$
@@
For 2-minute and 3-minute horizons, SWAG-SP achieves the best inMAE scores (0.54 and 0.77 minutes for 2 and 3 minutes), outperforming Bayesian [12] and IIA-Net [13]. At the 5-minute horizon, SWAG-SP ranks second with its wMAE score, showing strong adaptability for mid-range anticipations. IIA-Net [13] uses additional manual annotations like instrument bounding boxes and segmentation maps for supervision, while our approach relies solely on phase labels.
In Table 3, we compare our SWAG-SP method with previous approaches on the Remaining Surgery Duration (RSD) task. Using 4-fold cross-validation, our model ranks second on MAE-5 and MAE-ALL, outperforming previous methods, except BD-Net, all methods exclusively designed for this task. These results underscore our model’s versatility across several benchmarks.
@@
Fig. 4 illustrate a more practical representation of our approach, seamlessly extending the well-established phase recognition task with our generative-based anticipation framework. We show how the remaining time predictions per class (last eight rectangles) can be stacked into a 1-dimensional future segment (transparent) after the recognised surgical phases in an online manner (red dashed vertical line).
We show additional qualitative results and extended ablations in the supplementary materials (Sec. 1 in SM).
Figure 4: Surgical phase recognition and anticipation within a 30-minute horizon (y-axis) for Cholec80’s video 68, with the current elapsed time (x-axis) indicated by the red dashed vertical line. The thin, second row shows the ground truth phase segments, while the top rectangle displays the classified frames into phases; note that the more transparent colours to the right of the dashed line indicate the generated future phases at 60-second intervals. Each per-class regressed vertical segment is mapped onto a 1D timeline for both past and future. The last eight rectangles illustrate the regression-to-classification predictions per class.
@@ 5 Discussion
Our study evaluates methods for anticipating future surgical phases using transformer-based generative approaches. We investigate two decoding approaches: single-pass (SWAG-SP) and auto-regressive (SWAG-AR). We then demonstrate how prior knowledge can be used to initialise future tokens in single-pass decoding (SWAG-SP*), resulting in improved performance. Additionaly, we show how the remaining time prediction task with the single-pass model (SWAG-SP) can be extended to a classification task (SWAG-SP-R2C).
While the auto-regressive (SWAG-AR) model works well for short-term phase anticipation, its performance declines over longer horizons, accumulating errors over time. Our results show that SWAG-SP achieves the lowest wMAE and inMAE for up to 5-minute horizons, outperforming both the Bayesian model [12] and IIA-Net [13]. While the Bayesian model provides uncertainty estimation, it lacks short-term precision, and IIA-Net’s reliance on bounding boxes and segmentation masks may limit its scaling feasibility. We then extend the anticipation horizon beyond the previous 5-minute horizon to 15, 30, 45, and 60 minutes, bridging into the remaining surgery duration (RSD) prediction task, ranking second, while not making specific changes to the model.
Although our experiments indicate that using single-pass decoding with remaining time regression supervision achieves comparable quantitative performance to class-based generative approaches (SWAG-AR and SWAG-SP*), sequence generation offers the advantage of predicting repeated class occurrences within a single horizon, enabling the generation of fine-grained action sequences in a procedure. While the ability to predict multiple class occurrences within a sequence enhances model flexibility, it also increases the risk of misclassification on workflows with many class types and transitions. Cholecystectomies in Cholec80 follow a more standardised protocol with stronger phase ordering and fewer class repetitions [32], making transitions more predictable. In contrast, AutoLaparo21 exhibits greater variability in phase ordering and repetition [33]. While both datasets contain the same number of phase classes, these structural differences likely impact anticipation performance.
Real-time phase anticipation can alleviate the cognitive load on surgical teams and support intraoperative decision-making by offering foresight into upcoming phases over whole surgeries. This predictive support could improve team coordination, reduce surgery times, and enhance patient safety. Nevertheless, for wide clinical acceptance, future approaches must incorporate robust uncertainty estimation, as exemplified by the Bayesian perspective [12].
However, precise intraoperative workflow timing prediction faces intrinsic challenges due to anatomical differences, surgeon skill, patient-specific factors, and real-time changes. Surgery is not entirely deterministic, with multiple possible paths leading to optimal outcomes, making accurate long-term prediction especially challenging for generative models, especially with few training samples. We therefore distinguish two perspectives: (1) anticipating a surgeon’s future surgical workflow intraoperatively, which is highly variable over long horizons, offers limited accuracy even for high-level events like surgical phases, when trained on small datasets, as demonstrated by our results ($\simeq 50\%$), and (2) generating plausible future workflows derived from annotated surgical videos. Even if surgery is infinitely complex to predict in a time-sensitive manner, our SWAG approach could provide valuable guidance in structured scenarios. Moving forward, we believe that future methods should adopt a more open-ended, generative perspective, accommodating multiple plausible surgical workflow trajectories that can all provide optimal outcomes for the patient rather than assuming a single valid path.
@@ 6 Conclusion
In this work, we introduced SWAG, a novel model unifying surgical phase recognition and long-term generative-based anticipation within a single encoder-decoder framework enhanced by successfully embedding prior knowledge in future-generated tokens. Evaluations on the Cholec80 and AutoLaparo21 datasets demonstrate that SWAG outperforms naive baselines in classification and achieves competitive results in regression tasks, specifically for remaining surgery duration predictions. To support its clinical integration, we have demonstrated SWAG’s utility within a software application [^1] https://maxboels.github.io/swag. This work establishes SWAG as a novel framework in surgical workflow anticipation to support intraoperative guidance.

### v3

@@ 3.3 Experimental Setup
Datasets. We evaluate our work on two publicly available surgical phase datasets: Cholec80 (C80) [32] and AutoLaparo21 (AL21) [33]. For Cholec80, we adopt the same dataset splits as in [34], using 32 videos for training, 8 for validation, and 40 for testing. Similarly, for AL21, we apply the 10/4/7 split for training, validation, and testing, following [33, 5]. For the regression task, we follow a 60/20 train-test split as in [12, 13]. Both datasets were sampled at 1 frame per second (fps) following previous works [20, 4]. Across both datasets, we train on the 7 surgical phases, including an additional end-of-surgery (EOS) class, while disregarding tool annotations.
Training Details. We train our model on a maximal horizon of $N=h_{N}$ minutes and evaluate it at intermediate steps without re-training. We use a simple weighted cross-entropy loss for the classification task and mean squared error loss for the regression task. Additional details on the training procedure are provided in the supplementary material (Sec. 1.1 in SM).
Anticipation Time. Our method evaluates phase anticipation across multiple time horizons to assess both short-term and long-term predictive capabilities. This approach is justified by the substantial duration of procedures in our datasets: C80 training procedures average $38$ minutes, while AL21 training videos average $66$ minutes. By supervising our models on extended time horizons that encompass most of these workflows, we gain an important practical benefit: the ability to estimate the completion time of the entire surgical procedure from its earliest stages.
Classification Task. Since long-term future phase classification has not been studied in previous work, direct comparison to other surgical workflow anticipation methods is not possible. Therefore, we assess our generative approaches against two baseline methods: a simple continuation model (Naive1) and a conditional-probabilistic model (Naive2). Naive1 baseline extends the current recognised class over the future predicted horizon for $N$ minutes. While effective for short-term predictions, its performance naturally degrades over time with new class transitions. Naive2 baseline predicts future phases by sampling from the class probability distribution $P_{i,j,h_{n}\cdot 60}$, conditioned on the predicted current class $i$ and at the anticipated time index $h_{n}$. This baseline is effective due to the strong priors inherent in structured workflows, leveraging high recognition accuracy and precisely defined anticipation times.
@@ 3.4 Evaluation Metrics
We use a combination of classification and regression metrics to evaluate our methodology. These metrics are calculated for each anticipation time, allowing us to analyse the model’s performance through time. For classification, we use weighted F1 scores which account for class imbalance by assigning importance proportional to class frequency. Surgical workflows are naturally imbalanced due to the inherent structure of procedural steps. To limit EOS class dominance, we cap EOS samples at 4 minutes for Cholec80 and 8 minutes for AutoLaparo21. We also introduce SegF1, a segment-based F1 score that evaluates temporal coherence by matching predicted segments to ground truth using IoU thresholding, penalizing oversegmentation while rewarding correct phase boundaries. Detailed calculation is provided in the Supplementary Material. For regression tasks, we use the Mean Absolute Error (MAE) to evaluate the model’s performance on predicting the remaining time until the next phase occurrence. This metric quantifies the average absolute prediction errors in minutes. We use the wMAE, inMAE, and outMAE for weighting samples outside and inside the temporal horizon [12, 13]. Model selection is performed separately for each dataset using the validation set. For classification, we select the model with the highest overall mean weighted F1 score. For regression, we choose the model with the lowest weighted MAE.
@@ 4.1 Surgical Phase Anticipation
Table 1 compares our models against two naive baselines. While Naive2 achieves strong anticipation performance on Cholec80 (F1 = 39.5%), our SP* model excels on the more challenging AutoLaparo21 dataset (F1 = 41.3%), demonstrating superior generalisation to complex, variable procedures. Notably, despite Naive2’s strong minute-level results, it performs poorly on multiple minutes intervals due to discontinuous predictions, whereas our approach maintains consistent segment-level (SegF1) performance. Figure 3 illustrates performance degradation across anticipation horizons.
Table 1: Surgical Phase Recognition, Anticipation, and Segment-level Performance. Frame-level metrics use weighted averaging to handle class imbalance. Recognition: accuracy/F1-score at current time ($t=0$). Anticipation: mean F1-score over horizon and IoU-based segment F1.
Methods Cholec80 AutoLaparo21
Recognition Anticipation Recognition Anticipation
Acc F1 F1 SegF1 Acc F1 F1 SegF1
Naive1 90.8 91.1 29.3 24.5 70.7 71.0 19.6 16.4
Naive2 90.8 91.2 39.5 11.9 69.4 69.3 34.3 10.7
AR 90.3 90.7 27.8 25.0 73.7 72.9 29.3 23.3
R2C 89.6 89.8 36.1 32.5 72.4 71.8 32.9 29.2
SP 88.3 88.8 29.4 23.8 74.6 73.5 38.4 30.8
SP 88.3 88.8 32.1 29.8 73.3 72.7 41.3 34.8
  • Note: Frame-level metrics use weighted averaging for class imbalance. Recognition: Acc/F1 at $t=0$. Anticipation: mean F1 over horizon. Segment: IoU-based F1 with EOS weighting (max 4/8 EOS samples per sequence). Best performance highlighted, second-best underlined. Baseline models (recognition only). Prior knowledge initialization.
Figure 3: Models performance for surgical phase recognition and anticipation on Cholec80 and AutoLaparo21. Accuracy ($Acc_{[1:h_{n}]}$) over $h_{n}$ minutes (top right corners).
Figures 4 and 5 provide qualitative comparisons for recognition and anticipation tasks across both datasets. Top-performing cases show predicted phase sequences that closely match ground truth timing and durations, while bottom-performing videos exhibit systematic temporal misalignment where predictions consistently over- or under-estimate phase lengths. The model maintains smooth, continuous predictions across all performance levels.
Figure 4: Qualitative results on high- and mid-performing test samples for phase recognition and anticipation using the SWAG-SP* model. Each block represents a single video from either Cholec80 or AutoLaparo21: the top two rows correspond to top-performing cases, and the bottom two to median-performing ones. For each video, the upper panel shows predicted phase segments over a 30-minute anticipation horizon; the lower panel shows the corresponding ground-truth. The $x$-axis indicates the current time within the surgery; the $y$-axis corresponds to the future anticipation horizon. Colours represent surgical phases, with grey denoting the end-of-surgery class.
Figure 5: Qualitative results on two bottom-performing test samples using the SWAG-SP* model. Each example shows predicted (top) and ground-truth (bottom) phase segments over a 30-minute anticipation horizon. The $x$-axis represents surgical time progression, and the $y$-axis denotes anticipation into the future. Colours indicate surgical phases; grey corresponds to the end-of-surgery class.
@@ 4.2 Surgical Phase Anticipation with Remaining Time
Based on the classification evaluation results, we select the single-pass (SWAG-SP) decoder for comparison on state-of-the-art methods for remaining time until the next phase occurrence (see Table 2). For 2-minute and 3-minute horizons, SWAG-SP achieves the best inMAE scores (0.54 and 0.77 minutes for 2 and 3 minutes), outperforming Bayesian [12] and IIA-Net [13]. At the 5-minute horizon, SWAG-SP ranks second with its wMAE score, showing strong adaptability for mid-range anticipations.
@@
wMAE []delimited-[][$\downarrow$][ ↓ ] inMAE []delimited-[][$\downarrow$][ ↓ ] outMAE []delimited-[][$\downarrow$][ ↓ ]
@@
IIA-Net [13] uses additional manual annotations like instrument bounding boxes and segmentation maps for supervision, while our approach relies solely on phase labels. In Table 3, we compare our SWAG-SP method with previous approaches on the Remaining Surgery Duration (RSD) task. Using 4-fold cross-validation, our model ranks second on MAE-5 and MAE-ALL, outperforming previous methods, except BD-Net, all methods exclusively designed for this task.
@@
Fig.  6 illustrates how predicted future phase segments would be depicted during inference. We extend the standard visualisation method for the recognition task by completing the missing part on the right side in an online setting with the predicted remaining surgical workflow. Our method proposed to fill this gap in a simple continuous approach using dense temporal segment classification, similarly to the recognition task but with a forward thinking approach. We believe this novel segment-based anticipation visualisation method could improve intraoperative awareness and guidance.
Figure 6: Surgical phase recognition and anticipation within a 30-minute horizon (y-axis) for Cholec80’s video 68, with the current elapsed time (x-axis) indicated by the red dashed vertical line. The thin, second row shows the ground truth phase segments, while the top rectangle displays the classified frames into phases; note that the more transparent colours to the right of the dashed line indicate the generated future phases at 60-second intervals. The last eight rectangles depict the previous remaining time predictions per class mapping to the 1-d timeline at the top.
@@ 5 Discussion
Our study reformulates surgical phase anticipation as a generative sequence modeling task, introducing SWAG with transformer-based approaches under auto-regressive (AR) and single-pass (SP) decoding strategies, enhanced through token-level conditioning with prior knowledge (SP*).
Our results reveal distinct performance patterns that reflect the underlying complexity of different surgical procedures. On Cholec80, with its highly structured and predictable workflow, the Naive2 baseline achieves remarkably strong anticipation performance (F1 = 39.5%), demonstrating that simple approaches can excel when procedures follow consistent phase ordering. However, this advantage disappears on the more complex AutoLaparo21 dataset, where our SP* model demonstrates clear superiority (F1 = 41.3%, SegF1 = 34.8%), significantly outperforming naive baselines. This contrast highlights a fundamental insight: as surgical complexity increases, sophisticated modeling approaches become essential for accurate anticipation. The differential performance of our approaches provides valuable insights into anticipation strategies. Single-pass decoding with remaining time regression (R2C) achieves stronger performance on structured procedures like Cholec80 (F1 = 36.1% vs 32.9% on AutoLaparo21), suggesting that regression-based approaches work well when phase durations are predictable. Conversely, sequence generation approaches (SP*) excel on variable procedures, better capturing the complex temporal dependencies and phase transitions characteristic of less standardised surgeries. The superior segment-level performance of our approaches compared to naive methods (despite similar frame-level scores) demonstrates the importance of maintaining temporal coherence in predictions. We successfully extend anticipation horizons beyond the previous 5-minute limit, with SWAG-SP achieving the lowest wMAE and inMAE for short-term horizons while bridging into remaining surgery duration prediction. However, performance degrades significantly beyond 15-20 minute horizons, with F1 scores dropping below 30% for most methods, indicating fundamental challenges in long-term surgical workflow prediction even for standardized procedures.
Limitations and Future Directions
Our work reveals that long-term surgical anticipation remains challenging, with performance rapidly degrading over extended horizons due to the inherent unpredictability of intraoperative events and decision-making. Current limitations include reliance on relatively small datasets and the assumption of single valid future trajectories. Future work should prioritise three key directions: (1) incorporating uncertainty estimation to better handle the stochastic nature of surgical workflows, (2) developing evaluation frameworks that accommodate multiple plausible surgical trajectories rather than assuming deterministic workflows, and (3) scaling to larger, more diverse surgical datasets enriched with fine-grained action annotations and instrument control data. Looking toward longer-term advances, exploring generative approaches to create novel surgical scenarios beyond existing expert demonstrations could train robots to handle unprecedented cases, enhancing the current standard of surgical care and elevating patient safety to unprecedented levels.
@@ 6 Conclusion
This work introduces SWAG, a unified encoder-decoder framework for surgical phase recognition and long-term anticipation, successfully demonstrating the value of generative sequence modeling enhanced with prior knowledge conditioning. Our evaluation across Cholec80 and AutoLaparo21 datasets reveals that while simple approaches suffice for structured procedures, sophisticated modeling becomes essential for complex, variable surgeries. SWAG’s superior performance on challenging datasets and its ability to maintain temporal coherence in predictions establishes it as a promising foundation for intraoperative guidance systems, while highlighting the fundamental challenges that remain in long-term surgical workflow anticipation.